OpenStudy (anonymous):

Find the product of the complex number and its conjugate. 1 + 2i

OpenStudy (aravindg):

Let's take a+ib as the complex number. Then conjugate is a-ib. Can you multiply them and see?

OpenStudy (anonymous):

(a+ib)(a-ib) a^2-aib aib -2ib a^2-2ib

OpenStudy (aravindg):

ib*ib=i^2b^2

OpenStudy (aravindg):

Now i^2=-1 so u get -b^2

OpenStudy (aravindg):

So now u have a^2-(-b^2)=a^2+b^2